OpenStudy (atianaaaa):

could anyone solve 1/16=64^4x-3 ? without looking it up

OpenStudy (kingogcobraxx):

much do 4x-3 first very simple x=7/12

OpenStudy (kingogcobraxx):

must not much oops

OpenStudy (atianaaaa):

i know the answer was just looking for the work

geerky42 (geerky42):

\[\dfrac1{16}=64^{4x-3}\]right?

geerky42 (geerky42):

\[\dfrac1{16}=64^{4x-3}\\~\\4^{-2}=(4^3)^{4x-3}\\~\\4^{-2}=4^{12x-9}\\~\\-2=12x-9\\~\\7 = 12x\\~\\\boxed{x=\dfrac{7}{12}}\]

OpenStudy (atianaaaa):

hey can you help me with another one ? @geerky42